Given isosceles trapezoid ABCD, find the following.
AC (in cm), if BD = 15.2 cm
AC = ?cm
x, if AC = 3(x + 7)and BD = 8(x − 6)
x =

Answer by mananth(16946)



The diagonals of an isosceles trapezium are equal
BD =AC =15.2 CM
if AC = 3(x + 7)and BD = 8(x − 6)
3(x + 7)= 8(x − 6)
3x+21 = 8x-48
rearrange
8x-3x = 21+48
5x = 69
x = 13.8